# Refinement of Monocular Depth Maps via Multi-View Differentiable Rendering

### v1

@@ Appendix C Further Ablation studies
Setting the initial downsample factor to $d$ = 4 (reducing 1738k values in the depth map to 224k vertices) showed consistently better results than $d$ = 8, as shown in Tab. 6.
Mesh decimation ratios $r$ of 1.0, 0.5 (reducing to 112k vertices), and 0.25 (56k vertices) show very negligible differences, with 0.5 slightly outperforming 1.0 on some metrics. Based on this, we chose $r$= 0.5 as the default.
Setup RMSE MAE L1-rel L1-inv Acc. Acc. Acc.
0.01 m 0.05 m 0.10 m
@@
Comparing different monocular depth estimation techniques (see Tab. 7), both Marigold and Depth Anything Large (L) perform nearly equally well. This suggests our method might benefit from future advancements in monocular estimators. Furthermore, although the Depth Anything Small (S) model produces results in seconds compared to minutes for Marigold, its performance is only marginally lower than its larger counterparts. Based on the overall slight lead, we chose Marigold as the default for our method.
Coordinate Refinement.
Fig. 8 shows the influence of the $(u,v)$ coordinate refinement, as introduced in Main-Sec. 3.7, thus counterbalancing the intial resolution loss by setting $d$ = 4. This refinement significantly improved sharpness and accuracy in our depth maps, see Main-Tab. 3.
Figure 8: Refined edges through $(u,v)$ coordinate refinement. Left crops: Initial depth map is exhibits jaggy edges along silhouettes. Right crops: After convergence, silhouettes are piece-wise linearly approximated by triangle edges.
Intermediate Results.
Method RMSE MAE L1-rel L1-inv $\delta_{1}$ Acc. Acc. Acc.
0.01 m 0.05 m 0.10 m
Init Scale 0.17 0.11 0.09 0.10 0.93 0.10 0.41 0.65
After Local Ref. 0.16 0.09 0.07 0.10 0.93 0.21 0.58 0.76
Table 8: Depth map quality evaluated on our Scannet++ test scenes. Especially the accuracy metrics highlight the improvement gained through the local refinement (e.g. 110 % improvement in case of $\tau$=0.01 m, while MAE was only lowered by $\approx$20 %).
Tab. 8 shows the results after our intial global scaling and after the local refinement.
We can see great results, especially in the absolute, closeness based metric through our refinement strategies.

### v2

@@ Appendix F Ablation Studies
In Main-Sec. 3.3, introduced the intial downample factor $d$ for the triangulation and the quadric mesh decimation ratio $r$. In Main-Tab. 1, we show that the meshing step, where dd=4 was used, has an negligible impact on the depth map. Using $d$ = 4 reduced the 1738k values in the depth map to 224k vertices (87 % reduction). Additionally, we evaluated with $d$ = 8 and different mesh decimation ratios $r$, as shown in Tab. 10.
Setting the initial downsample factor to $d$ = 4 showed consistently better results than $d$ = 8. Mesh decimation ratios $r$ of 1.0, 0.5 (reducing to 112k vertices), and 0.25 (56k vertices) show very negligible differences, with 0.5 slightly outperforming 1.0 on some metrics. Based on this, we chose rr= 0.5 as the default.
Setup RMSE MAE L1-rel L1-inv $\uparrow Acc.$ Acc.
$\downarrow$ 0.01 m 0.05 m 0.10 m
@@
Coordinate Refinement.
Fig. 12 shows the influence of the $(u,v)$ coordinate refinement, as introduced in Main-Sec. 3.6, thus counterbalancing the initial resolution loss by setting $d$ = 4. This refinement significantly improved sharpness and accuracy in our depth maps, see Tab. 6.
Figure 12: Refined edges through $(u,v)$ coordinate refinement. Left crops: Initial depth map exhibits jaggy edges along silhouettes. Right crops: After convergence, silhouettes are piece-wise linearly approximated by triangle edges.
@@
Comparing different monocular depth estimation techniques (see Tab. 11), both Marigold and Depth Anything2 Large (L) perform nearly equally well. This suggests our method might benefit from future advancements in monocular estimators. Furthermore, although the DepthAnything2 Small (S) model produces results in seconds compared to minutes for Marigold, its performance is only marginally lower than its larger counterparts. In case of DepthAnything2 metric, we did not apply our global scale estimation and instead relied on the absolute values of the depth map. The qualitative differences indicate that metric prediction was still comparably inaccurate which could not be outweighed by later optimization steps.. Based on the overall slight lead, we chose Marigold as the default for our method if not indicated otherwise.
Tonemapping.
Figure 13: Tonemapping. In datasets with high variance in exposures (such as Tanks and Temples Ignatius), our tonemapping module still allows photometric optimization.
Tonemapping improves results for datasets like Ignatius which exhibit high variances of camera parameters like exposure or response. In this case, precision went up from 0.52 to 0.55, recall from 0.77 to 0.86, F-score from 0.62 to 0.67.
An example of the impact on the input images is shown in Fig. 13. Such balanced images could also be used to produce more evenly colored point clouds.
Dense Geometric Consistency & Multi Depth Map Refinement.
Moreover, we explored extending our technique to concurrently optimizing multiple depth maps. To do so, we selected additional views of the image set $I_{n}$ to predict depth maps. Then, we follow our global estimation and meshing procedure for additional views of the image set $I_{n}$ to get intial depth maps. During their concurrent refinement, we 1) add a geometric consistency loss $\mathcal{L}_{\textit{gc}}$} that enforces that projected depth values that have similar color also should have equal depth and 2) employ a simple novel view synthesis algorithm that does alpha composition of the warped depth maps to better resolve visibility (instead of masking) to calculate another photometric loss $\mathcal{L}_{\textit{photo}}^{\text{MV}}$.
As shown in Tab. 12, qualitative differences are minor and we could not determine noteworthy benefits. However, the number of optimizable parameters scales with the number of depth maps and the introduced losses expand the computational graph for backpropagation further. Consequently, the required VRAM increases drastically (possible batch size had to be reduced from 14 to 4 in case of 24 GB VRAM), so we decided against further usage.
Setup RMSE MAE L1-rel L1-inv $\delta_{1}$ Acc. Acc. Acc.
$\downarrow$ $\uparrow$ 0.01 m 0.05 m 0.10 m
$\scriptstyle\mathcal{L}_{\textit{photo}}$ 0.194 0.091 0.095 0.112 0.906 0.262 0.605 0.760
$\scriptstyle\mathcal{L}_{\textit{photo}}^{\text{MV}}$ 0.175 0.088 0.093 0.112 0.920 0.193 0.584 0.777
$\scriptstyle{\mathcal{L}_{\textit{gc}}}$ 0.179 0.091 0.097 0.111 0.920 0.188 0.578 0.773
$\scriptstyle\mathcal{L}_{\textit{photo}}+{\mathcal{L}_{\textit{gc}}}+\mathcal{L}_{\textit{photo}}^{\text{MV}}$ 0.193 0.093 0.094 0.129 0.895 0.260 0.593 0.745
Table 12: Ablation study for a multiview setup.